OpenStudy (anonymous):

can someone explain how sec(-pi/4) = sqrt2? I have never encountered a negative angle like this roflmao

OpenStudy (anonymous):

sec(-pi/4) = 1/(cos(-pi/4)) which equals sqrt of 2

OpenStudy (blockcolder):

You just answered your question. =))

OpenStudy (anonymous):

yeah but I don't know why or how this is looking at the unit circle

OpenStudy (anonymous):

cosine deals with values on the x-axis, sin deals with values on the y-axis.  notice how at that angle, even though y is negative, the x value is still positive? so when you do: 1/cos( -pi/4) you will still get a positive value. since cos (-pi/4)=cos(pi/4)= sqrt(2)/2 1/cos(-pi/4)= 2/sqrt(2)=sqrt(2)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

conjugation x_x